# Healing Powers of BERT: How Task-Specific Fine-Tuning Recovers Corrupted Language Models

### v1

@@ Abstract
Language models like BERT excel at sentence classification tasks due to extensive pre-training on general data, but their robustness to parameter corruption is unexplored. To understand this better, we look at what happens if a language model is "broken", in the sense that some of its parameters are corrupted and then recovered by fine-tuning. Strategically corrupting BERT variants at different levels, we find corrupted models struggle to fully recover their original performance, with higher corruption causing more severe degradation. Notably, bottom-layer corruption affecting fundamental linguistic features is more detrimental than top-layer corruption. Our insights contribute to understanding language model robustness and adaptability under adverse conditions, informing strategies for developing resilient NLP systems against parameter perturbations.
Healing Powers of BERT: How Task-Specific Fine-Tuning Recovers Corrupted Language Models
Shijie Han 11footnotemark: 1 Columbia University sh4460@columbia.edu                        Zhenyu Zhang thanks: Equal contribution Zhejiang University zhenyuzhang@zju.edu.cn
Andrei Arsene Simion Columbia University aas2148@columbia.edu
@@ 1 Introduction
Pre-trained language models like BERT Kenton and Toutanova (2019) have propelled tremendous progress in sentence classification, an important natural language processing (NLP) task. The powerful representational capabilities of BERT enable more accurate and robust classification across diverse domains Yin et al. (2019). While BERT demonstrates exceptional ability in capturing intricate linguistic nuances, enabling it to analyze entire sentences and discern their meanings and relationships Reimers and Gurevych (2019), the inner workings of BERT are still not fully understood.
Several studies have investigated the robustness of BERT to various perturbations Dong et al. (2021); Hauser et al. (2023), but the effects of parameter corruption on fine-tuning performance for specific tasks remain unexplored. In this paper, we explore the impact of controlled parameter corruption on BERT’s ability to perform sentence classification tasks. We hypothesize that corrupting a pre-trained BERT model will hinder its fine-tuned performance on sentence classification tasks, and then we explore how BERT recovers its original performance during task-specific fine-tuning.
@@ 2 Related Work
BERT (Bidirectional Encoder Representations from Transformers) revolutionized NLP with its pre-training and fine-tuning methods Kenton and Toutanova (2019). Unlike GPT-style unidirectional models, BERT uses a bidirectional transformer architecture, understanding word context from both directions. This bidirectional approach and its rich token representations enable BERT to excel in various natural language understanding tasks.
@@
BERT pre-training involves two objectives: Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). In MLM, some input tokens are randomly masked, and the model learns to predict the masked tokens using the surrounding context. This allows BERT to capture nuanced language patterns and dependencies by considering context from both sides of a token Radford et al. . The NSP task further enhances BERT by training it to predict if a given sentence B logically follows sentence A Liu et al. (2019); Sanh (2019). NSP helps model understand relationships between sentences, which is important for tasks involving sentence pairs Peters et al. (2018).
@@
Once pre-trained, BERT can be fine-tuned for specific downstream tasks with minimal architectural modifications. For text classification, a simple linear layer is added on top of the pre-trained model, using the output of the special [CLS] token as the aggregate sequence representation Kenton and Toutanova (2019); Sun et al. (2019). During fine-tuning, the entire model is trained on labeled data for the target task. BERT’s rich pre-trained knowledge provides a strong foundation that can accurately classify text based on nuanced linguistic features when fine-tuned on labeled datasets Yin et al. (2019). The [CLS] token output is passed through an additional classification layer to produce the final output, making BERT highly versatile for various text classification applications like sentiment analysis and topic categorization Wang et al. (2018); Hoang et al. (2019); Xu et al. (2019); George and Sumathy (2023).
@@
Despite the impressive performance of the BERT model, its inherent complex architecture also makes it vulnerable to various forms of corruption, such as adversarial attacks and weight perturbations, which can severely impact its performance Ebrahimi et al. (2018). While some research has attempted targeted interventions to influence model performance based on an understanding of the Transformer architecture, our understanding of how its internal components react to corruption remains limited Meng et al. (2022); Dai et al. (2022); Geva et al. (2021); De Cao et al. (2021). This highlights the need for a deeper understanding of the model, as enhancing our comprehension of BERT’s robustness is crucial for model interpretability and reliable deployment Feder et al. (2021). To bridge this gap, we conduct an in-depth investigation to reveal the extent to which BERT can recover through fine-tuning after being corrupted.
@@ 3 Corruption Method
The real-world corruption of BERT is likely to be random and unpredictable, rather than following the same distribution as the pre-trained parameters. Therefore, we simulate corruption in BERT by randomly initializing their internal parameters. Specifically, for corrupting the parameters of the LayerNorm layers, we standardize all weight parameters by setting them to a constant value of 1.0. For the weights of non-LayerNorm layers, we re-initialize them using the Kaiming uniform initialization method He et al. (2015), which draws samples from a uniform distribution with specific bounds determined by the input size of the layer. Additionally, all bias parameters are directly initialized to zero.
@@ 4.1 Experimental Settings
To assess the robustness of language models for parameter corruption during task-specific fine-tuning, we choose three BERT models with different scales: BERT-Base with 110M parameters Kenton and Toutanova (2019), BERT-Large with 340M parameters Kenton and Toutanova (2019), and DistilBERT with 66M parameters Sanh (2019). We corrupt these models using the method outlined in Section 3 and then fine-tune and evaluate them on six classification datasets spanning sentiment analysis and topic categorization: SST-2 Socher et al. (2013), IMBD Maas et al. (2011), AG News Zhang et al. (2015), Emotion Saravia et al. (2018), DBPedia Zhang et al. (2015) and Twitter Financial News Topic (TFNT) Magic (2022). To control our experiments, we balance the learning difficulty on a dataset with the available computational resources by truncating portions of the dataset and redefining the training and testing splits, as detailed in Table 2.
Dateset Categories Train Test Balanced
SST-2 2 56k 12k Yes
IMDB 2 25k 5k Yes
AG News 4 48k 6k Yes
Emotion 6 14k 4k No
DBPedia 14 50k 12k Yes
TFNT 20 17k 4k No
Table 2: Classification datasets details
@@
Our implementation utilizes the following software and hardware configurations: PyTorch Paszke et al. (2019) version 1.13.1; Transformers Wolf et al. (2020) library version 4.41.1; CUDA version 12.4; GPU: one NVIDIA RTX4090D-24G; We employ different learning rates of 1e-5, 2e-5, and 5e-5 and choose the best, training over 10 epochs for each dataset with Adam Kingma and Ba (2015) as our optimizer.
4.2 Results
@@
4.3 Feature Similarity.
We employ the t-SNE Van der Maaten and Hinton (2008) algorithm to project [CLS] tokens gained from the fine-tuned BERT models with various corruption patterns to 2D space. t-SNE visualization reveals clusters in features, which help in understanding the classification performance and observing intrinsic feature patterns and relationships.
4.4 Outlier Analysis.
From our results, intriguing outliers emerge such as a low accuracy of 33.97% for BERT-Large on IMDB with 50% bottom corruption, and 16.98% accuracy on Emotion with 25% and 50% bottom corruption. These anomalies, marked by convergence failures and low metrics, may result from the sensitivity of BERT-Large’s middle layers, the complexity of its larger parameter set, and specific characteristics of the IMDB and Emotion datasets. These factors highlight the complex interplay between datasets, architectures, and corruption patterns, sometimes leading to unexpected outcomes.
5 Conclusion
Limitation
Here are a few limitations of this work:
(1) Corruption Method: Our study employs a specific method of parameter corruption by random initialization. Other types of corruption, such as noise injection or targeted parameter manipulation, can potentially yield different insights.
(2) Model Architectures: The research focuses on the BERT family of language models. The impact of parameter corruption may vary across different model architectures, such as recurrent neural networks, convolutional models, or other Transformer-based models.
(3) Fine-tuning Strategies: The study explores the standard fine-tuning approach, where all model parameters are updated during training. Alternative fine-tuning strategies, such as layer-wise adaptation or low-rank adaptation, could potentially mitigate or exacerbate the effects of parameter corruption.
Appendix
Appendix A More Analysis
To further verify the rules summarized in the main paper, we use different experimental settings to perform the following analysis:
@@ Appendix A More Analysis
A.1 Average Token
A.2 Linear Probe
In previous experiments, we fine-tune all BERT model parameters to assess the impact of parameter corruption on performance. Alternatively, we can use a linear probing approach by freezing the pre-trained BERT parameters and only fine-tuning the classification head. This method allows us to investigate how well corrupted BERT models transfer learned representations to downstream tasks without further adapting the main model parameters.
The results presented in Table 4 appear highly unstable and inconsistent, making it challenging to observe clear trends identified when the entire BERT is fine-tuned. One potential reason for this instability is the optimization process for the classification head may struggle to converge, leading to erratic behavior and performance fluctuations. These irregularities and inconsistencies suggest that the linear probing approach may not be suited for accurately assessing the impact of parameter corruption on the model’s performance. The lack of adaptation in the main BERT body could exacerbate the effects of corruption, leading to unstable and unpredictable results.
A.3 Non-classification Task
@@ Appendix B License of Scientific Artifacts
There is no specific license. Anyone using the dataset should cite the Maas et al. (2011)
@@
There is no specific license in this page. Anyone using the dataset should cite the Ando and Zhang (2005).
@@
The Recognizing Textual Entailment (RTE) datasets come from a series of annual textual entailment challenges. The authors of the GLUE benchmark combined the data from RTE1 Dagan et al. (2006), RTE2 Bar Haim et al. (2006), RTE3 Giampiccolo et al. (2007) and RTE5 Bentivogli et al. (2009). Examples are constructed based on news and Wikipedia text. Anyone using the dataset should cite above and GLUE Wang et al. (2019). There is no specific license in tac page and acl page.

### v2

@@ Abstract
Language models like BERT excel at sentence classification tasks due to extensive pre-training on general data, but their robustness to parameter corruption is unexplored. To understand this better, we look at what happens if a language model is ”broken”, in the sense that some of its parameters are corrupted and then recovered by fine-tuning. Strategically corrupting BERT variants at different levels, we find corrupted models struggle to fully recover their original performance, with higher corruption causing more severe degradation. Notably, bottom-layer corruption affecting fundamental linguistic features is more detrimental than top-layer corruption. Our insights contribute to understanding language model robustness and adaptability under adverse conditions, informing strategies for developing resilient NLP systems against parameter perturbations.
Robustness, Hierarchical & Concept explanations, Probing
@@ 1 Introduction
Pre-trained language models like BERT (Kenton & Toutanova, 2019) have propelled tremendous progress in sentence classification, an important natural language processing (NLP) task. The powerful representational capabilities of BERT enable more accurate and robust classification across diverse domains (Yin et al., 2019). While BERT demonstrates exceptional ability in capturing intricate linguistic nuances, enabling it to analyze entire sentences and discern their meanings and relationships (Reimers & Gurevych, 2019), the inner workings of BERT are still not fully understood.
Several studies have investigated the robustness of BERT to various perturbations (Dong et al., 2021; Hauser et al., 2023), but the effects of parameter corruption on fine-tuning performance for specific tasks remain unexplored. In this paper, we explore the impact of controlled parameter corruption on BERT’s ability to perform sentence classification tasks. We hypothesize that corrupting a pre-trained BERT model will hinder its fine-tuned performance on sentence classification tasks, and then we explore how BERT recovers its original performance during task-specific fine-tuning.
@@ 2 Related Work
BERT (Bidirectional Encoder Representations from Transformers) revolutionized NLP with its pre-training and fine-tuning methods (Kenton & Toutanova, 2019). Unlike GPT-style unidirectional models, BERT uses a bidirectional transformer architecture, understanding word context from both directions. This bidirectional approach and its rich token representations enable BERT to excel in various natural language understanding tasks.
@@
BERT pre-training involves two objectives: Masked Language Modeling (MLM) and Next Sentence Prediction (NSP). In MLM, some input tokens are randomly masked, and the model learns to predict the masked tokens using the surrounding context. This allows BERT to capture nuanced language patterns and dependencies by considering context from both sides of a token (Radford et al., ). The NSP task further enhances BERT by training it to predict if a given sentence B logically follows sentence A (Liu et al., 2019; Sanh, 2019). NSP helps model understand relationships between sentences, which is important for tasks involving sentence pairs (Peters et al., 2018).
@@
Once pre-trained, BERT can be fine-tuned for specific downstream tasks with minimal architectural modifications. For text classification, a simple linear layer is added on top of the pre-trained model, using the output of the special [CLS] token as the aggregate sequence representation (Kenton & Toutanova, 2019; Sun et al., 2019). During fine-tuning, the entire model is trained on labeled data for the target task. BERT’s rich pre-trained knowledge provides a strong foundation that can accurately classify text based on nuanced linguistic features when fine-tuned on labeled datasets (Yin et al., 2019). The [CLS] token output is passed through an additional classification layer to produce the final output, making BERT highly versatile for various text classification applications like sentiment analysis and topic categorization (Wang et al., 2018; Hoang et al., 2019; Xu et al., 2019; George & Sumathy, 2023).
@@
Despite the impressive performance of the BERT model, its inherent complex architecture also makes it vulnerable to various forms of corruption, such as adversarial attacks and weight perturbations, which can severely impact its performance (Ebrahimi et al., 2018). While some research has attempted targeted interventions to influence model performance based on an understanding of the Transformer architecture, our understanding of how its internal components react to corruption remains limited (Meng et al., 2022; Dai et al., 2022; Geva et al., 2021; De Cao et al., 2021). This highlights the need for a deeper understanding of the model, as enhancing our comprehension of BERT’s robustness is crucial for model interpretability and reliable deployment (Feder et al., 2021). To bridge this gap, we conduct an in-depth investigation to reveal the extent to which BERT can recover through fine-tuning after being corrupted.
@@ 3 Corruption Method
The real-world corruption of BERT is likely to be random and unpredictable, rather than following the same distribution as the pre-trained parameters. Therefore, we simulate corruption in BERT by randomly initializing their internal parameters. Specifically, for corrupting the parameters of the LayerNorm layers, we standardize all weight parameters by setting them to a constant value of 1.0. For the weights of non-LayerNorm layers, we re-initialize them using the Kaiming uniform initialization method (He et al., 2015), which draws samples from a uniform distribution with specific bounds determined by the input size of the layer. Additionally, all bias parameters are directly initialized to zero.
@@
Table 1: Classification datasets details
Dateset Categories Train Test Balanced
SST-2 2 56k 12k Yes
IMDB 2 25k 5k Yes
AG News 4 48k 6k Yes
Emotion 6 14k 4k No
DBPedia 14 50k 12k Yes
TFNT 20 17k 4k No
@@ 4.1 Experimental Settings
To assess the robustness of language models for parameter corruption during task-specific fine-tuning, we choose three BERT models with different scales: BERT-Base with 110M parameters (Kenton & Toutanova, 2019), BERT-Large with 340M parameters (Kenton & Toutanova, 2019), and DistilBERT with 66M parameters (Sanh, 2019). We corrupt these models using the method outlined in Section 3 and then fine-tune and evaluate them on six classification datasets spanning sentiment analysis and topic categorization: SST-2 (Socher et al., 2013), IMBD (Maas et al., 2011), AG News (Zhang et al., 2015), Emotion (Saravia et al., 2018), DBPedia (Zhang et al., 2015) and Twitter Financial News Topic (TFNT) (Magic, 2022). To control our experiments, we balance the learning difficulty on a dataset with the available computational resources by truncating portions of the dataset and redefining the training and testing splits, as detailed in Table 1.
@@
Our implementation utilizes the following software and hardware configurations: PyTorch (Paszke et al., 2019) version 1.13.1; Transformers (Wolf et al., 2020) library version 4.41.1; CUDA version 12.4; GPU: one NVIDIA RTX4090D-24G; We employ different learning rates of 1e-5, 2e-5, and 5e-5 and choose the best, training over 10 epochs for each dataset with Adam (Kingma & Ba, 2015) as our optimizer.
4.2 Main Results
@@
4.3 Further Exploration
Average token instead of CLS token.
@@ 4.3 Further Exploration
Linear Probe.
We have fine-tuned all BERT model parameters to assess the impact of parameter corruption on performance. Alternatively, we can use a linear probing approach by freezing the pre-trained BERT parameters and only fine-tuning the classification head. This method allows us to investigate how well corrupted BERT models transfer learned representations to downstream tasks without further adapting the main model parameters. The results presented in Table 4 appear highly unstable and inconsistent, making it challenging to observe clear trends identified when the entire BERT is fine-tuned. One potential reason for this instability is that the optimization process for the classification head may struggle to converge, leading to erratic behavior and performance fluctuations. These irregularities and inconsistencies suggest that the linear probing approach may not be suited for accurately assessing the impact of parameter corruption on the model’s performance. The lack of adaptation in the main BERT body could exacerbate the effects of corruption, leading to unstable and unpredictable results.
Non-classification Tasks.
@@
4.4 Feature Similarity
We employ the t-SNE (Van der Maaten & Hinton, 2008) algorithm to project [CLS] tokens gained from the fine-tuned BERT models with various corruption patterns to 2D space. t-SNE visualization reveals clusters in features, which help in understanding the classification performance and observing intrinsic feature patterns and relationships.
5 Conclusion
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A More Analysis: Outlier Analysis.
From our results in Table 2, intriguing outliers emerge, such as a low accuracy of 33.97% for BERT-Large on IMDB with 50% bottom corruption, and 16.98% accuracy on Emotion with 25% and 50% bottom corruption. These anomalies, marked by convergence failures and low metrics, may result from the sensitivity of BERT-Large’s middle layers, the complexity of its larger parameter set, and specific characteristics of the IMDB and Emotion datasets. These factors highlight the complex interplay between datasets, architectures, and corruption patterns, sometimes leading to unexpected outcomes.
@@ Appendix B License of Scientific Artifacts
There is no specific license. Anyone using the dataset should cite the (Maas et al., 2011)
@@
There is no specific license in this page. Anyone using the dataset should cite the (Ando & Zhang, 2005).
@@
The Recognizing Textual Entailment (RTE) datasets come from a series of annual textual entailment challenges. The authors of the GLUE benchmark combined the data from RTE1 (Dagan et al., 2006), RTE2 (Bar Haim et al., 2006), RTE3 (Giampiccolo et al., 2007) and RTE5 (Bentivogli et al., 2009). Examples are constructed based on news and Wikipedia text. Anyone using the dataset should cite above and GLUE (Wang et al., 2019). There is no specific license in tac page and acl page.
Appendix C Limitation
Here are a few limitations of this work:
  • Corruption Method: Our study employs a specific method of parameter corruption by random initialization. Other types of corruption, such as noise injection or targeted parameter manipulation, can potentially yield different insights.
  • Model Architectures: The research focuses on the BERT family of language models. The impact of parameter corruption may vary across different model architectures, such as recurrent neural networks, convolutional models, or other Transformer-based models.
  • Fine-tuning Strategies: The study explores the standard fine-tuning approach, where all model parameters are updated during training. Alternative fine-tuning strategies, such as layer-wise adaptation or low-rank adaptation, could potentially mitigate or exacerbate the effects of parameter corruption.